OpenStudy (anonymous):

You are trying to determine the half-life of a new radioactive element you have isolated. You start with 1 gram, and 4 days later you determine that it has decayed down to 0.6 grams. What is its half-life? (Round your answer to three significant digits.)

OpenStudy (dumbcow):

start with general exponential decay function: \[P = e^{-kt}\] plug in given point (4,0.6) solve for "k" \[k = \frac{\ln 0.6}{-4}\] to determine half-life, set P = 0.5 , solve for "t" \[t = \frac{4* \ln 0.5}{\ln 0.6}\]